OpenStudy (adajiamcneal):

The application rate of a liquid lawn fertilizer is 2 ounces of fertilizer per 100 sq feet of lawn are. The fertilizer is only sole in whole quarts (1 quart = 32 oz) What is the smallest number of quarts of fertilizer that can be purchased to cover a lawn with an area of 4960 sq feet

OpenStudy (cj49):

m not sure but 4960/100=49.6*2=99.2 99.2/32=3.1

OpenStudy (anonymous):

Ceiling[4960/1600]=4 Ceiling gives the smallest integer greater than or equal to x.

OpenStudy (anonymous):

1600 = ( 32/2 )*100

OpenStudy (adajiamcneal):

thank you